# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v5

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution-free framework for constructing well-calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a split-conformal calibration layer on a rolling window and, in its TCP-RM variant, augments the conformal threshold with a single online Robbins-Monro (RM) offset to steer coverage toward a target level in real time. We benchmark TCP against GARCH, Historical Simulation, a rolling tree-based Quantile Regression (QR) model, a classical linear quantile regression baseline (QR-Linear), and an adaptive conformal method (ACI) across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold). Three results are consistent across assets. First, both QR and QR-Linear yield the sharpest intervals but are materially under-calibrated, and even ACI remains below the nominal 95% target in our full-sample backtests. Second, TCP (and TCP-RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs. 5.06). Third, the RM update changes calibration and width only marginally at our default hyperparameters. Crisis-window visualizations around March 2020 show TCP/TCP-RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage). A sensitivity study confirms robustness to window size and step-size choices. Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@ 1.  INTRODUCTION
Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz1952; Morgan1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion2007; Dowd1998), assumptions that can systematically understate tail risk. Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar2000). Events like the 2008 global financial crisis (Brunnermeier2009) and the March 2020 stock-market crash (Mazur2021) exposed the fragility of parametric assumptions, especially in the tails.
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often exhibit heavy tails (Cont2001). GARCH-type models (Engle1982; Bollerslev1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk2005). However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies. Several works extend conformal methods to sequential or temporally dependent settings (Xu2021; Stankeviciute2021; Gibbs2021), but a robust and practical solution for financial markets remains an open challenge.
@@ 2.  RELATED WORK
Traditional Financial Risk Models. Quantitative risk management traces its roots to portfolio theory (Markowitz1952) and was formalized around the VaR concept with J.P. Morgan’s RiskMetrics (Morgan1996). In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion2007; Dowd1998; Glasserman2004). Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar2000). However, many traditional approaches underperform in crisis backtests (Kupiec1995; Christoffersen1998), highlighting a blind spot when tail events dominate.
Advanced Time-Series and Machine-Learning Models. To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle1982; Bollerslev1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson1991; Glosten1993). For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (Pickands1975; McNeil2000), and VaR under heavy tails raises specific concerns (Danielsson2000). On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker1978), with ensemble and neural variants such as quantile regression forests (Meinshausen2006) and quantile networks (Taylor2000); deep LSTMs capture complex temporal dependencies (Fischer2018). We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke2017).
Conformal Prediction Theory. Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk2005). Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs2021; Xu2021; Stankeviciute2021). Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
@@
Recent research has addressed non-stationary environments by adapting the conformal guarantee. Gibbs2021 and zaffran2022adaptive introduce Adaptive Conformal Inference (ACI), which updates the target miscoverage rate ($\alpha_{t}$) online based on past coverage errors. While effective, we argue that in heavy-tailed financial settings, oscillating the target probability can lead to unstable interval widths; TCP-RM instead adapts the conformal threshold ($C_{t}$) directly via a Robbins-Monro update, offering a stable, scalar lever for width control. Other approaches, such as CPTD (lin2022temporal), leverage cross-sectional exchangeability across a panel of time series to restore validity. In contrast, TCP is designed for the univariate setting (single asset) where cross-sectional pooling (e.g., mixing Bitcoin and Gold) is inappropriate due to distinct, idiosyncratic asset dynamics.
@@ 4.  MATHEMATICAL BACKGROUND
This section lays the theoretical groundwork for our proposed method. Our calibration mechanism is related in spirit to adaptive conformal procedures for nonstationary settings (zaffran2022adaptive; lin2022temporal) but differs in using a Robbins–Monro update on the conformal threshold while holding the nominal level fixed. We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
@@ 4.3 Temporal Conformal Prediction (TCP)
    Stochastic approximation framework. The Robbins-Monro update for the offset is analyzed using classical stochastic-approximation arguments under mixing and bounded-variance noise; our assumptions instantiate those conditions in this setting ((Kushner2003)).
@@
    Assumption verification in practice. We empirically probe the mixing/independence assumptions using standard risk diagnostics (Kupiec1995 and Christoffersen1998 backtests); the lack of serial dependence in exceedances supports the validity of these conditions on our financial datasets.
@@ 5.2.2 Benchmark Models
    Adaptive Conformal Inference (ACI): A rolling implementation of Gibbs2021 that uses the same window split as TCP ($w_{\mathrm{tr}}=192$ training, $w_{\mathrm{cal}}=60$ calibration). However, instead of updating the conformal threshold $C_{t}$ (like TCP-RM), ACI updates the target miscoverage rate $\alpha_{t}$ online based on the most recent coverage error (step size $\eta=0.05$).
@@ 5.3 Evaluation Framework
For each model and asset, we also perform standard 1-day-ahead VaR backtests on the 5% lower tail. Specifically, we report: (i) the unconditional coverage test of Kupiec1995 (LRuc), (ii) the independence test of Christoffersen1998 (LRind), and (iii) their sum $_{\mathrm{cc}}{=}LR$$_{\mathrm{uc}}{+}$LRind for conditional coverage. Each test yields a likelihood-ratio statistic and a $p$-value, together with the number of exceedances $x$ (days where the realized return falls below the nominal 5% VaR) out of $n$ out-of-sample predictions.
@@ 6.1 Backtests of tail coverage
To complement marginal coverage and interval width, we perform standard VaR-style backtests on the 5% lower tail using the Kupiec (Kupiec1995) unconditional coverage test (LR_uc), the Christoffersen (Christoffersen1998) independence test (LR_ind), and their joint conditional coverage test (LR_cc). Table 3 reports the likelihood-ratio statistics, $p$-values, the total number of exceedances (days where the realized return falls below the nominal 5% lower bound), and the number of out-of-sample predictions $n$ for the S&P 500.
@@
We also perform sensitivity analyses for TCP-RM and TCP, with results reported in Appendix F and Appendix G. TCP vs. TCP-RM trace plots are presented in Appendix H.
@@ 7.  CONCLUSION AND FUTURE WORK
Code Availability
@@ Appendix B ADDITIONAL CRISIS-WINDOW RESULTS
Appendix C ADDITIONAL TCP VS TCP-RM CRISIS-WINDOW PLOTS
@@ Appendix F SENSITIVITY ANALYSIS FOR TCP-RM
This section reports the full hyperparameter sensitivity analysis for TCP-RM across the S&P 500, BTC-USD, and Gold. For each asset we vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$. For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Tables 7,8, and 9).
Hyperparameters Performance
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9575 5.3052
100 0.010 0.9575 5.3066
100 0.050 0.9581 5.3172
252 0.005 0.9579 5.1168
252 0.010 0.9579 5.1191
252 0.050 0.9586 5.1314
500 0.005 0.9525 5.4135
500 0.010 0.9525 5.4319
500 0.050 0.9567 5.5348
@@
Summary. On S&P 500, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), with w=100w{=}100 and w=500w{=}500 producing slightly wider bands and coverage close to or just above the 95% target, except for mild undercoverage in some w=500w{=}500 settings; this motivates our choice w=252w{=}252, $\gamma_{0}{=}0.01$ in the main experiments. For BTC-USD, interval widths shrink as $w$ increases, with the tightest bands at w=500w{=}500, while coverage remains stably around 95–96% across all configurations. For Gold, widths are smallest at w=100w{=}100 and grow with $w$, and coverage stays near the nominal level throughout. Across all three assets, TCP-RM is therefore stable over the $(w,\gamma_{0})$ grid considered here: the window length $w$ controls the sharpness–stability tradeoff, whereas the specific choice of $\gamma_{0}$ within this range has only a minor effect on coverage and width.
Appendix G SENSITIVITY ANALYSIS ON TCP (NO RM)

### v6

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution-free framework for constructing well-calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a rolling split-conformal calibration layer; its TCP-RM variant adds an online Robbins-Monro offset to steer coverage in real time. We benchmark TCP against GARCH, Historical Simulation, Quantile Regression (QR), linear QR, and Adaptive Conformal Inference (ACI) across S&P 500, Bitcoin, and Gold. Three results are consistent. First, QR baselines yield the sharpest intervals but are materially under-calibrated; even ACI remains below the 95% target. Second, TCP achieves near-nominal coverage, yielding intervals slightly wider than Historical Simulation (e.g., S&P 500: 5.21 vs. 5.06). Third, the RM update changes calibration only marginally at default hyperparameters. Crisis-window visualizations (March 2020) show TCP promptly expanding and contracting intervals as volatility spikes. A sensitivity study confirms robustness to hyperparameters. Overall, TCP bridges statistical inference and machine learning, providing a practical solution for calibrated risk forecasting under distribution shift.
@@ 1.  Introduction
Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz, 1952; J.P. Morgan/Reuters, 1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion, 2007; Dowd, 1998), assumptions that can systematically understate tail risk. Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar and Uryasev, 2000). Events like the 2008 global financial crisis (Brunnermeier, 2009) and the March 2020 stock-market crash (Mazur et al., 2021) exposed the fragility of parametric assumptions, especially in the tails.
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often exhibit heavy tails (Cont, 2001). GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005). However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies. Several works extend conformal methods to sequential or temporally dependent settings (Xu and Xie, 2021; Stankeviciute et al., 2021; Gibbs and Candès, 2021), but a robust and practical solution for financial markets remains an open challenge.
@@ 2.  Related Work
Traditional Financial Risk Models. Quantitative risk management traces its roots to portfolio theory (Markowitz, 1952) and was formalized around the VaR concept with J.P. Morgan’s RiskMetrics (J.P. Morgan/Reuters, 1996). In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion, 2007; Dowd, 1998; Glasserman, 2004). Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar and Uryasev, 2000). However, many traditional approaches underperform in crisis backtests (Kupiec, 1995; Christoffersen, 1998), highlighting a blind spot when tail events dominate.
Advanced Time-Series and Machine-Learning Models. To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle, 1982; Bollerslev, 1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson, 1991; Glosten et al., 1993). For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (Pickands, 1975; McNeil and Frey, 2000), and VaR under heavy tails raises specific concerns (Danielsson and de Vries, 2000). On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker and Bassett, 1978), with ensemble and neural variants such as quantile regression forests (Meinshausen, 2006) and quantile networks (Taylor, 2000); deep LSTMs capture complex temporal dependencies (Fischer and Krauss, 2018). We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke et al., 2017).
Conformal Prediction Theory. Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk et al., 2005). Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs and Candès, 2021; Xu and Xie, 2021; Stankeviciute et al., 2021). Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
@@
Recent research has addressed non-stationary environments by adapting the conformal guarantee. Gibbs and Candès (2021) and Zaffran et al. (2022) introduce Adaptive Conformal Inference (ACI), which updates the target miscoverage rate ($\alpha_{t}$) online based on past coverage errors. While effective, we argue that in heavy-tailed financial settings, oscillating the target probability can lead to unstable interval widths; TCP-RM instead adapts the conformal threshold ($C_{t}$) directly via a Robbins-Monro update, offering a stable, scalar lever for width control. Other approaches, such as CPTD (Lin et al., 2022), leverage cross-sectional exchangeability across a panel of time series to restore validity. In contrast, TCP is designed for the univariate setting (single asset) where cross-sectional pooling (e.g., mixing Bitcoin and Gold) is inappropriate due to distinct, idiosyncratic asset dynamics.
@@ 4.  Mathematical Background
This section lays the theoretical groundwork for our proposed method. Our calibration mechanism is related in spirit to adaptive conformal procedures for nonstationary settings (Zaffran et al., 2022; Lin et al., 2022) but differs in using a Robbins–Monro update on the conformal threshold while holding the nominal level fixed. We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
@@ 4.3 Temporal Conformal Prediction (TCP)
    Stochastic approximation framework. The Robbins-Monro update for the offset is analyzed using classical stochastic-approximation arguments under mixing and bounded-variance noise; our assumptions instantiate those conditions in this setting ((Kushner and Yin, 2003)).
@@
    Assumption verification in practice. We empirically probe the mixing/independence assumptions using standard risk diagnostics (Kupiec (1995) and Christoffersen (1998) backtests); the lack of serial dependence in exceedances supports the validity of these conditions on our financial datasets.
@@ 5.2.2 Benchmark Models
    Adaptive Conformal Inference (ACI): A rolling implementation of Gibbs and Candès (2021) that uses the same window split as TCP ($w_{\mathrm{tr}}=192$ training, $w_{\mathrm{cal}}=60$ calibration). However, instead of updating the conformal threshold $C_{t}$ (like TCP-RM), ACI updates the target miscoverage rate $\alpha_{t}$ online based on the most recent coverage error (step size $\eta=0.05$).
@@ 5.3 Evaluation Framework
For each model and asset, we also perform standard 1-day-ahead VaR backtests on the 5% lower tail. Specifically, we report: (i) the unconditional coverage test of Kupiec (1995) (LRuc), (ii) the independence test of Christoffersen (1998) (LRind), and (iii) their sum $_{\mathrm{cc}}{=}LR$$_{\mathrm{uc}}{+}$LRind for conditional coverage. Each test yields a likelihood-ratio statistic and a $p$-value, together with the number of exceedances $x$ (days where the realized return falls below the nominal 5% VaR) out of $n$ out-of-sample predictions.
@@ 6.1 Backtests of tail coverage
To complement marginal coverage and interval width, we perform standard VaR-style backtests on the 5% lower tail using the Kupiec (Kupiec, 1995) unconditional coverage test (LR_uc), the Christoffersen (Christoffersen, 1998) independence test (LR_ind), and their joint conditional coverage test (LR_cc). Table 3 reports the likelihood-ratio statistics, $p$-values, the total number of exceedances (days where the realized return falls below the nominal 5% lower bound), and the number of out-of-sample predictions $n$ for the S&P 500.
@@
6.2 Sensitivity Analysis
This section reports the full hyperparameter sensitivity analysis for TCP-RM across the S&P. We vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$. For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Table 4).
Hyperparameters Performance
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9575 5.3052
100 0.010 0.9575 5.3066
100 0.050 0.9581 5.3172
252 0.005 0.9579 5.1168
252 0.010 0.9579 5.1191
252 0.050 0.9586 5.1314
500 0.005 0.9525 5.4135
500 0.010 0.9525 5.4319
500 0.050 0.9567 5.5348
On S&P 500, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), with w=100w{=}100 and w=500w{=}500 producing slightly wider bands and coverage close to or just above the 95% target, except for mild undercoverage in some w=500w{=}500 settings; this motivates our choice w=252w{=}252, $\gamma_{0}{=}0.01$ in the main experiments.
We also perform sensitivity analyses for TCP-RM on BTC-USD and Gold and TCP (on all three assets), with results reported in Appendix F and Appendix G.
TCP vs. TCP-RM trace plots are presented in Appendix H.
@@ 7.  Conclusion and Future Work
Code and Data Availability
The complete implementation of the proposed Temporal Conformal Prediction (TCP) framework is publicly available at https://github.com/agnivibes/temporal-conformal-prediction-tcp/tree/main. The repository contains all source code required to reproduce the results reported in this paper, along with the processed financial returns dataset (financial_returns.csv) used in the empirical analysis.
@@ Appendix B Additional Crisis-window Results
Appendix C Additional TCP vs TCP-RM Crisis-window Results
@@ Appendix F Sensitivity Analysis for TCP-RM
This section reports the full hyperparameter sensitivity analysis for TCP-RM across BTC-USD, and Gold. For each asset we vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$. For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Tables 8, and 9).
@@
Summary. For BTC-USD, interval widths shrink as $w$ increases, with the tightest bands at w=500w{=}500, while coverage remains stably around 95–96% across all configurations. For Gold, widths are smallest at w=100w{=}100 and grow with $w$, and coverage stays near the nominal level throughout. Across all three assets, TCP-RM is therefore stable over the $(w,\gamma_{0})$ grid considered here: the window length $w$ controls the sharpness–stability tradeoff, whereas the specific choice of $\gamma_{0}$ within this range has only a minor effect on coverage and width.
Appendix G Sensitivity Analysis for TCP (no RM)